### v1

# Scalable learning of potentials to predict time-dependent Hartree-Fock dynamics

Abstract
We propose a framework to learn the time-dependent Hartree-Fock (TDHF) inter-electronic potential of a molecule from its electron density dynamics. Though the entire TDHF Hamiltonian, including the inter-electronic potential, can be computed from first principles, we use this problem as a testbed to develop strategies that can be applied to learn a priori unknown terms that arise in other methods/approaches to quantum dynamics, e.g., emerging problems such as learning exchange-correlation potentials for time-dependent density functional theory. We develop, train, and test three models of the TDHF inter-electronic potential, each parameterized by a four-index tensor of size up to $60\times 60\times 60\times 60$. Two of the models preserve Hermitian symmetry, while one model preserves an eight-fold permutation symmetry that implies Hermitian symmetry. Across seven different molecular systems, we find that accounting for the deeper eight-fold symmetry leads to the best-performing model across three metrics: training efficiency, test set predictive power, and direct comparison of true and learned inter-electronic potentials. All three models, when trained on ensembles of field-free trajectories, generate accurate electron dynamics predictions even in a field-on regime that lies outside the training set. To enable our models to scale to large molecular systems, we derive expressions for Jacobian-vector products that enable iterative, matrix-free training.
I Introduction
To understand phenomena such as charge transfer, spectroscopy, and the response of a molecule to an applied laser field, it is essential to be able to compute the dynamics of electrons. For such systems, the governing equation is the time-dependent electronic Schrödinger equation (TDSE):
$\imath\frac{d\Psi(\mathbf{r},t)}{dt}=\widehat{H}(\mathbf{r},t)\Psi(\mathbf{r},t).$ (1)
Here $\imath=\sqrt{-1}$ denotes the imaginary unit, and $\widehat{H}(\mathbf{r},t)$ is the electronic Hamiltonian operator that operates on the time-dependent many-body electronic wave function $\Psi(\mathbf{r},t)$, where $\mathbf{r}$ represents the spatial and spin coordinates of all electrons. The word electronic signifies that other particles (i.e., protons and more generally, nuclei) are held fixed as part of the Born-Oppenheimer approximation, and we solve for the electronic wave function in the field of fixed nuclei. For real, multi-electron molecular systems of interest, direct, grid-based numerical solution of (1) is not possible. Additionally, for many properties of interest, the full wave function $\Psi$ is unnecessary; the $1$-electron reduced density $\rho$, a much lower-dimensional quantity (itself computable from $\Psi$), would suffice.
In principle, time-dependent density functional theory (TDDFT) enables one to compute the reduced electron density $\rho$, without first solving the TDSE for $\Psi$ (Ullrich, 2012). However, in TDDFT, the density-dependent exchange-correlation potential (itself a part of the overall inter-electronic potential) is not known and must be approximated (Gross and Maitra, 2012; Maitra, 2016). TDDFT is widely used to simulate electron dynamics and excited state properties, but in some cases the approximations employed for the exchange-correlation potential lead to inaccurate electron dynamics and charge transfer (Elliott et al., 2012; Rappoport and Hutter, 2012; Habenicht et al., 2014; Provorse, Habenicht, and Isborn, 2015; Lacombe and Maitra, 2020; Ranka and Isborn, 2023). Learning a more accurate Hamiltonian, including the unknown density-dependent electron potential of TDDFT, would lead to advances in the simulation of the molecular properties and excited state dynamics.
At the time of writing, there are only three works we know of that seek to learn the TDDFT exchange-correlation potential from data consisting of temporal snapshots of the electron density (Suzuki, Nagai, and Haruyama, 2020; Bhat et al., 2022; Yang and Whitfield, 2023). These works focus on one-dimensional model problems; extending the methods to three-dimensional systems of chemical/physical interest will be a large-scale undertaking from both a theoretical and computational point of view. We lack an understanding of questions such as: (i) how should we form our training sets? (ii) how should we construct our model potentials? and (iii) how should we quantify errors? The literature on learning potentials for quantum molecular systems is tilted heavily toward static, time-independent problems (Snyder et al., 2012; Bartók et al., 2017; Chandrasekaran et al., 2019; Schleder et al., 2019; Jørgensen et al., 2019; Smith et al., 2019; Ceriotti, 2019) and offers little guidance on the answers to these questions, especially for molecular systems with, e.g., dozens of electrons.
The broad goal of this paper is to extend our understanding of how to learn potentials from time series observations of quantum systems. We develop a learning framework for dynamics governed by the time-dependent Hartree-Fock (TDHF) equation. We focus on TDHF for four reasons. First, for TDHF, we know a priori how to compute the entire Hamiltonian, including the potential term we seek to learn. Hence we can compare our learned potentials against the ground truth. Second, TDHF is numerically tractable even for large molecular systems, so there is no a priori obstacle that stops us from generating training data. Third, Hartree-Fock theory incorporates an exact treatment of electron exchange. Consequently, the inter-electronic potential is purely a function of a $2$-electron integral tensor $(ij|kl)$—see (5)—that enjoys an $8$-fold permutation symmetry. Though existing density functionals that approximate electron exchange based on the local density do not have this symmetry, we expect an exact exchange-correlation functional to have symmetries similar to what we see in Hartree-Fock theory. This motivates studying how to incorporate symmetries into a machine learning model of a potential. Fourth, Hartree-Fock is itself the first model (approximating $\Psi$ using a single Slater determinant) in a sequence of increasingly accurate models that culminates in full configuration interaction (in which $\Psi$ is approximated using a linear combination of all possible Slater determinants). Time-dependent full configuration interaction yields numerical solutions of (1) that are exact within a finite-dimensional basis set (McWeeny, 1989).
Therefore, by studying and solving a machine learning problem for TDHF, we aim to produce insights that can be abstracted and generalized to learning problems that involve more accurate models for quantum dynamics. The present study aims to guide future efforts to learn a priori unknown terms that play important roles in time-dependent equations of motion. This includes, but is not limited to, learning the exchange-correlation potential for TDDFT (Suzuki, Nagai, and Haruyama, 2020; Bhat et al., 2022; Yang and Whitfield, 2023), closure terms for time-dependent reduced density matrix functional theory and similar theories (Blum, 2012; Akbari et al., 2012; Ferré, Filatov, and Huix-Rotllant, 2016; Bhat et al., 2024), and surrogate models for the Kadanoff-Baym equations (Bassi et al., 2023).
The TDHF equation
$\imath\frac{dP(t)}{dt}=\big{[}H(P(t),t),P(t)\big{]},$ (2)
can be derived from the TDSE by approximating $\Psi$ as a single Slater determinant computed from a finite set of spin orbitals (Szabo and Ostlund, 2012). Here $P$ is the $1$-electron reduced density matrix corresponding to $\Psi$, while the Hamiltonian matrix $H$ is the operator $\widehat{H}$ from (1) expressed in a finite-dimensional basis (McLachlan and Ball, 1964; McWeeny, 1989). Note that (2) only holds if one expresses $P$ and $H$ in orthonormal bases. The full, time-dependent Hamiltonian in (2) is
$H(P(t),t)=H(P(t))+V_{\text{ext}}(t).$ (3)
On the right-hand side, the field-free Hamiltonian $H(P(t))$ includes the kinetic energy of the electrons, the potential due to interaction with the fixed nuclei, and the density dependent inter-electronic repulsion term, also known as the Fock matrix in Hartree-Fock theory. In (3), $V_{\text{ext}}(t)$ is an optional, external, time-dependent potential, e.g., an applied electric field. In our prior work (Bhat, Ranka, and Isborn, 2020; Gupta et al., 2022), we initiated the project of learning a model $\widetilde{H}(P)$ of the true field-free Hamiltonian $H(P)$ from time series observations $\{P(t_{j})\}$ of electron density matrices.
Throughout the remainder of this paper, if we write $H(\cdot,\cdot)$ with two arguments, then we are referring to the full, time-dependent Hamiltonian, the left-hand side of (3). If we write $H(\cdot)$ with only one argument, as in $H(P(t))$ or $H(P)$, then we are referring to the field-free Hamiltonian.
In the present work, we move beyond prior work in three ways. First, we pay close attention to scalability to systems such as paranitroaniline ($\text{C}_{6}\text{H}_{6}\text{N}_{2}\text{O}_{2}$) in the STO-3G basis set, where the density matrices $P(t)$ are of size $60\times 60$. For $\text{C}_{6}\text{H}_{6}\text{N}_{2}\text{O}_{2}$, the methods developed in this paper enable learning an inter-electronic potential parameterized by a tensor with more than $1.6$ million entries, with which we can propagate electron dynamics with an overall error on the order of $10^{-7}$—see Figure 4. This is six orders of magnitude better than what we can achieve using methods from our prior work. We are unaware of other methods that can scale to this system size. In the present work, we propose three models and derive expressions for Jacobian-vector products that enable matrix-free, iterative training. We formulate our models in such a way that they preserve symmetries using linear algebra. The combination of these techniques leads to scalability.
Second, in the present work, we explore the generation of training sets that are both highly accurate and highly informative for machine learning purposes. We employ a higher-order numerical method that, for the largest molecular systems we study, avoids spurious oscillations. Using this method, we construct a training set consisting of an ensemble of field-free trajectories. The ensemble explores more of phase space, passing near more minima, maxima, and saddle points that characterize the true Hamiltonian. As we will show, models trained on this ensemble have superior predictive power as measured by the ability to generate field-on trajectories $\widetilde{P}(t)$ that match true field-on trajectories $P(t)$.
Third, at the core of the present work lies the idea of building physics—in the form of symmetries—into our machine learning models. Two of the models we propose preserve Hermitian symmetry of the Hamiltonian $\widetilde{H}(P)$. Another preserves a deeper $8$-fold permutation symmetry that underpins the inter-electronic term in the true field-free Hamiltonian $H(P)$. We algorithmically construct a basis for the space of all $4$-index tensors that satisfy this symmetry. We express our model using this basis, thereby ensuring that the symmetry is preserved. Our methods can be generalized to other tensors, symmetries, and systems.
For the largest molecular system we study, we find that the $8$-fold symmetry-preserving ensemble-trained model dramatically outperforms all other models in terms of propagation error. For all molecular systems studied here, the $8$-fold symmetry-preserving model leads to the smallest difference between learned and true Hamiltonians. While incorporating known symmetries into the model clearly makes physical sense, our results show that it benefits the machine learning enterprise as well. The $8$-fold symmetry-preserving model uses approximately one-eighth as many parameters as other models we consider. For all systems, the $8$-fold symmetry-preserving model requires orders of magnitude fewer iterations to train. These findings have important consequences for future machine learning efforts that involve time-dependent quantum systems trained on time-dependent data.
II Methods
Our learning approach is guided by the following question: what training data and what type of model do we need in order to learn a TDHF inter-electronic potential that yields the smallest test set error?
II.1 Choice of initial condition and time-stepping method
In order to train and test our models, we must have on hand a reliable set of electron dynamics trajectories. We now describe how we generate this data.
For a particular molecule and basis set, we apply standard electronic structure methods to compute the ground truth field-free Hamiltonian $H(P)$ and an initial condition $P(0)$ with an applied delta-kick perturbation. Before doing anything further, we evolve the density matrix forward in time to $P(2\delta t)$ with time step $\delta t=8.268\times 10^{-2}$ a.u. We then relabel this density matrix $P(2\delta t)$ as $P(0)$ and use this as our initial condition for all subsequent modeling/testing. We use a modified version of the Gaussian electronic structure code (Frisch, Trucks, and et. al, 2018) to compute the Hamiltonian $H(P)$ and the matrices $P(0)$ through $P(2\delta t)$.
To understand why we relabel $P(2\delta t)$ as $P(0)$, note that in our past (Bhat, Ranka, and Isborn, 2020; Gupta et al., 2022) and present work, our learning procedures depend on first estimating $dP/dt$ from $P(t)$ and then fitting $idP/dt$ to $[H,P]$. That is, we minimize the residual between the left- and right-hand sides of (2). As expected given the nature of the Dirac delta, our numerical estimates of $dP/dt$ are much more accurate if we move past the time of the initial delta-kick perturbation. This improves all learning procedures we have tested.
Equipped with $H(P)$ and $P(0)$, our task is to numerically integrate (2) to produce a trajectory $\{P(t_{j})\}$ of electron density matrices on a temporal grid $t_{j}=j\Delta t$ for $j=0,\ldots,J$. In our prior work (Bhat, Ranka, and Isborn, 2020; Gupta et al., 2022), we employed the second-order modified midpoint unitary transformation (MMUT) method (Li et al., 2005a). In the present work, for reasons we now explain, we employ a fourth-order scheme (Casas and Iserles, 2006a) that we abbreviate as CI4. We detail both schemes in Supplementary Information (SI) Section V.1. Note that the CI4 method is based on a fourth-order Magnus expansion. In a comparative study of time-stepping methods for the time-dependent Kohn-Sham equations, a fourth-order Magnus integrator emerged as the clear winner (Gémez Pueyo et al., 2018).
Figure 1: Preliminary long-term propagation test using the exact Hamiltonian and two different propagation methods, the second-order MMUT method (red) and the fourth-order CI4 method (black). The bottom-most plot shows a zoomed in view of the real part of one density matrix element as a function of time, clearly showing that MMUT leads to large-amplitude spurious oscillations that are not present in the CI4 solution.
We present in Figure 1 the results of solving (2) for $\text{C}_{6}\text{H}_{6}\text{N}_{2}\text{O}_{2}$ in the STO-3G basis using both schemes, MMUT (red curves) and CI4 (black curves). We set $\Delta t=8.268\times 10^{-4}$ a.u. and propagate (2) for $200000$ steps; we use only the field-free Hamiltonian $H(P)$ and delta-kick initial condition $P(0)$ computed as described above, with no external field. The $P(t)$ matrix for this system is of size $60\times 60$. We find that both schemes yield comparable results for the first $20000$ time steps, and hence omit these from our plots.
[…]
To see this more clearly, in the bottom-most plot of Figure 1, we have zoomed in on the real part of the $P_{21,21}(t)$ trajectory between $50000$ and $50100$ time steps. The large-scale, high-frequency oscillations in the MMUT curve should be viewed in sharp contrast to the totally flat behavior of the CI4 solution. We find similar behavior for other systems of similar size, e.g., $\text{C}_{2}\text{H}_{4}$ in 6-31+$\text{G}^{*}$ (with $P(t)$ of size $46\times 46$) and LiH in 6-311++$\text{G}^{**}$ (with $P(t)$ of size $29\times 29$). Experiments with MMUT showed that decreasing $\Delta t$ by a factor of $10$ delays the onset of this divergence between trajectories, but at prohibitive computational cost to reach the same final integration time.
Therefore, for the remainder of this work, for each molecular system (choice of molecule plus basis set), we use the CI4 (Casas and Iserles, 2006a) scheme to solve (2). This includes generation of training and test data using either ground truth Hamiltonians $H(P)$ or learned Hamiltonians $\widetilde{H}(P)$.
II.2 Linear models
We begin with a set of atomic orbitals $\{\phi_{j}\}$. Each $\phi_{j}$ is a function of a spatial coordinate $\mathbf{r}\in\mathbb{R}^{3}$; each $\phi_{j}$ is an approximate solution of Schrödinger’s equation. The atomic orbitals (AO) form a basis used to express molecular orbitals, and so the collection $\{\phi_{j}\}$ is often referred to as a basis set. Hartree-Fock electronic structure theory yields an expression for the ground truth Hamiltonian $H^{\text{AO}}$ (in the AO basis), a matrix-valued function of the electron density matrix $P^{\text{AO}}$ (in the AO basis). For any complex quantity $Z$, let $\overline{Z}$ denote its complex conjugate. Then the $(i,j)$-th component of $H^{\text{AO}}$ is
$H^{\text{AO}}(P^{\text{AO}})_{ij}=H^{\text{core.AO}}_{ij}+\sum_{kl}\left[(ij|lk)-\frac{1}{2}(ik|lj)\right]P^{\text{AO}}_{kl},$ (4)
where $H^{\text{core.AO}}$ comprises the kinetic and electron-nuclear part of the Hamiltonian matrix (again, in the AO basis), and $(ij|kl)$ is the $2$-electron integral defined by
[…]
$P$ $=\mathcal{X}^{-1}P^{\text{AO}}\mathcal{X}^{-\dagger}$ (6a)
$H$ $=\mathcal{X}^{\dagger}H^{\text{AO}}\mathcal{X}.$ (6b)
Let us use the shorthand $\mathcal{E}$ for the following transformation of a critical tensor—the tensor in square brackets in the Hamiltonian (4)—into the CO basis:
$\mathcal{E}_{abcd}=\sum_{ijkl}\overline{\mathcal{X}_{ia}}\mathcal{X}_{jb}\left[(ij|lk)-\frac{1}{2}(ik|lj)\right]\mathcal{X}_{kc}\overline{\mathcal{X}_{ld}}.$ (7)
Then (6) and (4) together give us the ground truth Hamiltonian in the CO basis:
$H(P)_{ab}=H^{\text{core}}_{ab}+\sum_{cd}\mathcal{E}_{abcd}P_{cd}.$ (8)
In this paper, unless specified otherwise via ${}^{\text{AO}}$ superscripts, all density and Hamiltonian matrices ($P$ and $H$) are in the CO basis. We use the CO basis because it is orthonormal, and the TDHF equation (2) holds only in an orthonormal basis. The AO basis is not guaranteed to be orthonormal—in practice, it is not. Abstractly, we think of (8) as
$H(P)=H^{\text{core}}+H_{1}(P),$ (9)
where $H_{1}(P)$ is the Hermitian, density-dependent, inter-electronic potential. In this work, we are interested in methods that scale well to $N=60$ and beyond. At $N=60$, the Hermitian matrix $H^{\text{core}}$ is determined by $N^{2}=3600$ real numbers. The inter-electronic potential $H_{1}(P)$ is determined by the $4$-index tensor $\mathcal{E}$; accounting for symmetries that we describe further below, $\mathcal{E}$ is determined by just over $1.6\times 10^{6}$ real numbers. As high-dimensionality manifests most prominently in the $H_{1}(P)$ term, we focus on learning this inter-electronic potential in the present work. In prior work, we have found that we can learn the constant term $H^{\text{core}}$ very accurately (Bhat, Ranka, and Isborn, 2020; Gupta et al., 2022).
Purely for instructive/motivational purposes, we describe a basic linear model for $H(P)$. We use boldfaced lowercase $\mathbf{p}\in\mathbb{C}^{N^{2}}$ and $\mathbf{h}\in\mathbb{C}^{N^{2}}$ to denote vectorized or flattened versions of $P$ and $H$, respectively. Then a basic linear model is
$\widetilde{\mathbf{h}}=\mathbf{b}^{\text{core}}+\mathbf{p}\mathbf{b}.$ (10)
Here we have an intercept $\mathbf{b}^{\text{core}}$, a complex vector of shape $N^{2}\times 1$, and a complex coefficient matrix $\mathbf{b}$ of shape $N^{2}\times N^{2}$. A key problem with (10) is that after we reshape $\widetilde{\mathbf{h}}$ into an $N\times N$ matrix $\widetilde{H}$, we have no guarantee that $\widetilde{H}$ is Hermitian, as is required to obtain physically meaningful results.
To build a physically accurate model, we must ensure that $\widetilde{H}$ is Hermitian. One way to do this is to model only diagonal and upper-triangular elements of $\widetilde{H}$, and then use symmetry to fill in the lower-triangular elements. We followed this approach in our earlier work (Bhat, Ranka, and Isborn, 2020; Gupta et al., 2022). However, our prior approach of manually tracking diagonal and upper-triangular entries of $\widetilde{H}$ causes our prior methods to scale poorly for large $N$.
We use $\widetilde{H}_{1}(P)$ to denote a model of the density-dependent, inter-electronic potential in (9). We present three scalable $\widetilde{H}_{1}(P)$, in increasing order of mathematical complexity. All three are linear in the entries of $P$. All three preserve Hermitian symmetry. The third model preserves an additional $8$-fold permutation symmetry that we detail below.
All three $\widetilde{H}_{1}(P)$ models we present can be computed using standard numerical linear algebra operations, without manually looping over particular sets of indices as in our previous work. As numerical linear algebra is highly optimized on both CPU and GPU architectures, this improves scalability of the model for larger $N$.
II.2.1 Tied Regression
Observe that if one wants to convert a matrix $M$ into a Hermitian matrix, one can simply use $(M+M^{\dagger})/2$, i.e., the Hermitian part of $M$. This motivates our first model. With $P^{R}$ and $P^{I}$ denoting the real and imaginary parts of an electron density matrix $P$, we introduce the tied regression model:
$R_{kl}(P)$ $=H^{\text{core}}_{kl}+\sum_{i,j}P^{R}_{ij}\beta_{ijkl}$ (11a)
$Q_{kl}(P)$ $=\sum_{i,j}P^{I}_{ij}\beta_{ijkl}$ (11b)
$\widetilde{H}(P)$ $=\frac{1}{2}[R(P^{R})+R(P^{R})^{T}]+\frac{\imath}{2}[Q(P^{I})-Q(P^{I})^{T}].$ (11c)
Here $R$, $Q$, $H^{\text{core}}$, and $\bm{\beta}$ are all real; T denotes ordinary matrix transpose. The idea is that $R$ and $Q$ respectively model the real and imaginary parts of the Hamiltonian. As in the ground truth Hamiltonian, we allow $R$ and $Q$ to depend, respectively, on the real and imaginary parts of the density matrix $P$. We do not need an intercept term in $Q$ because, as noted above, the constant Hamiltonian term $H^{\text{core}}$ is real. Let $M=R+\imath Q$ be the presymmetrized model; the model (11c) then consists of the symmetrization of $M$, i.e., $\widetilde{H}=(M+M^{\dagger})/2$.
We refer to this as the tied regression model because the coefficient tensor $\bm{\beta}$ that multiplies $P^{R}$ in (11a) is also used to multiply $P^{I}$ in (11b). In preliminary work, we began with a model that had
$Q_{kl}(P)=\sum_{i,j}P^{I}_{ij}\gamma^{1}_{ijkl}$ (12)
instead of (11b); for this model, the task is to learn both $\bm{\beta}$ and $\bm{\gamma}$. Model (11) effectively ties the parameters $\bm{\beta}$ and $\bm{\gamma}$ to one another (via $\bm{\beta}=\bm{\gamma}$), reducing the number of real parameters from $2N^{4}$ to $N^{4}$, and thus reducing training times. In SI Section V.2.1, we show that there exists $\bm{\beta}$ such that the tied model (11) matches the ground truth Hamiltonian (in the CO basis). Thus there is nothing gained by considering a model in which (12) replaces (11b).
II.2.2 Regression with Hermitian Representation
Though the symmetrization approach described above is simple, it is also potentially wasteful. In (11a-11b), we build full $N\times N$ models of $R$ and $Q$, but in (11c), we only use the symmetric part of $R$ and the antisymmetric part of $Q$. This begs the question of whether we can model only those parts of $R$ and $Q$ to begin with. We especially seek to do this with linear/multilinear algebraic methods that lead to fast numerical implementations.
We begin by introducing a basis for the space of real symmetric matrices. Given such a matrix of size $N\times N$, including the diagonal, there are a total of $N(N+1)/2$ upper-triangular entries. To each such entry with index $(i,j)$, let us associate a symmetric matrix that is identically zero except for entries of $1$ at locations $(i,j)$ and $(j,i)$. In case $i=j$, the matrix will have only one $1$ at position $(i,i)$. In this way, to each integer $k$ from $1$ to $N(N+1)/2$, we associate an $N\times N$ matrix $B^{R}_{k}$. Stacking these matrices, we form a three-index tensor $B^{R}$ of size $N\times N\times N(N+1)/2$. By construction, these matrices form a basis for the space of real symmetric matrices; any such matrix $S$ can be written as
$S_{ij}=\sum_{k=1}^{N(N+1)/2}B^{R}_{ijk}v_{k}$ (13)
for some $\mathbf{v}\in\mathbb{R}^{N(N+1)/2}$. Continuing, suppose we wish to form a real, $4$-index tensor $\bm{\beta}$ such that $\beta_{a,b,c,d}=\beta_{a,b,d,c}$ for all $1\leq a,b,c,d\leq N$. For each fixed choice of $(a,b)$, $\beta_{a,b,:,:}$ is a real symmetric matrix that can be represented using $B^{R}$ as in (13). Thus $\bm{\beta}$ is determined by a $3$-index tensor $\mathbf{v}$ such that
$\beta_{abcd}=\sum_{k=1}^{N(N+1)/2}B^{R}_{cdk}v_{abk}.$ (14)
Following a similar construction, we can create a basis for the space of real antisymmetric matrices. Now we exclude the diagonal, so there are a total of $N(N-1)/2$ upper-triangular entries. To each such entry with index $(i,j)$, we associate an antisymmetric matrix that is identically zero except for entries of $1$ and $-1$ at locations $(i,j)$ and $(j,i)$, respectively. In this way, to each integer $k$ from $1$ to $N(N-1)/2$, we associate an $N\times N$ matrix $B^{I}_{k}$. Stacking these matrices, we form a three-index tensor $B^{I}$ of size $N\times N\times N(N-1)/2$.
Now suppose we wish to form a real, $4$-index tensor $\bm{\gamma}$ such that $\gamma_{a,b,c,d}=-\gamma_{a,b,d,c}$ for all $1\leq a,b,c,d\leq N$. Following the same logic as above, $\bm{\gamma}$ is determined by a $3$-index tensor $\mathbf{w}$ such that
$\gamma_{abcd}=\sum_{k=1}^{N(N-1)/2}B^{I}_{cdk}w_{abk}.$ (15)
We now use (14) and (15) to build our Hermitian representation model
$\widetilde{H}(P)_{kl}=H^{\text{core}}_{kl}+\sum_{i,j}P^{R}_{ij}\beta_{ijkl}+\imath\sum_{i,j}P^{I}_{ij}\gamma_{ijkl}.$ (16)
The tensors $B^{R}$ and $B^{I}$ are known and constant, depending only on the dimension $N$. Treating $H^{\text{core}}$ as known, the coefficients of this model amount to the entries of $\mathbf{v}$ and $\mathbf{w}$, of size $N\times N\times N(N+1)/2$ and $N\times N\times N(N-1)/2$, respectively. There are a total of $N^{4}$ entries, meaning this model has precisely the same size as our tied regression model (11).
This model does not tie $\bm{\beta}$ and $\bm{\gamma}$ together. Let :::: denote matrix-tensor contraction, so that
$\sum_{i,j}P^{R}_{ij}\beta_{ijkl}=[P^{R}:\bm{\beta}]_{kl}.$
Then we can rewrite (16) as $\widetilde{H}(P)=R+\imath Q$ with $R=H^{\text{core}}+P^{R}:\bm{\beta}$ and $Q=P^{I}:\bm{\gamma}$. By virtue of constructing $\bm{\beta}$ and $\bm{\gamma}$ using (14-15), we guarantee that $R$ and $Q$ are symmetric and antisymmetric, respectively. Thus $\widetilde{H}(P)$ is Hermitian. We do not need a final symmetrization step, as in (11c).
In SI Section V.2.2, we show that there exists $\bm{\beta}$, $\bm{\gamma}$ such that the Hermitian representation model (16) matches the ground truth Hamiltonian (in the CO basis).
II.2.3 Eight-Fold Symmetry-Preserving Regression
The Hermitian symmetry preserved by the first two models (11) and (16) can be viewed as a consequence of a deeper set of symmetries. Assuming real atomic orbitals $\{\phi_{j}\}$, Brailsford and Hall (1971) show that the $4$-index tensor $(ij|kl)$ defined in (5) satisfies the $8$-fold symmetry
$(ij|kl)=(ji|lk)=(kl|ij)=(lk|ji)\\ =(ji|kl)=(lk|ij)=(ij|lk)=(kl|ji).$ (17)
Based on these observations, let us attempt to model the $4$-index tensor $(ij|kl)$ itself. The first part of our procedure is analogous to how we formed bases for the spaces of real symmetric and antisymmetric matrices, respectively, in Section II.2.2.
Algorithm 1 Populating the sparse tensor $\mathscr{S}$ that enables representation of $4$-index tensors with $8$-fold symmetry
1:$m=0$ and empty $N\times N\times N\times N$ tensor $\mathscr{B}$
2:for $i,j,k,l=1,\dots,N$ do
3:     if $\mathscr{B}(i,j,k,l)=0$ then
4:         $m\leftarrow m+1$
5:         for $c=1,\ldots,8$ do
6:              $\mathscr{B}(\sigma^{c}(i,j,k,l))\leftarrow 1$
7:              append the $5$-tuple $(\sigma^{c}(i,j,k,l),m)$ to the index list of sparse tensor $\mathscr{S}$
8:         end for
9:     end if
10:end for
Here we form a basis $\mathscr{S}$ for the space of $N\times N\times N\times N$ tensors that satisfy the $8$-fold symmetry (17)—as shown in Brailsford and Hall (1971), this space has dimension
$n_{T}=\frac{1}{8}N(N+1)(N^{2}+N+2).$ (18)
To store this basis efficiently, we let $\mathscr{S}$ be a sparse tensor of dimension $N\times N\times N\times N\times n_{T}$, determined entirely by a list of indices where its value is $1$.
To populate $\mathscr{S}$, we follow Algorithm 1. For $m=1,2,\ldots,8$, let $\sigma^{m}(i,j,k,l)$ be the $m$-th permutation seen in (17). For instance, $\sigma^{4}(i,j,k,l)=(l,k,j,i)$. Let $\mathcal{O}(i,j,k,l)$ denote $\{\sigma^{m}(i,j,k,l)\mid m=1,\ldots,8\}$, i.e., the orbit of the $4$-tuple $(i,j,k,l)$ under the group $\{\sigma^{m}\}_{m=1}^{8}$. In Algorithm 1, each value of $m$ corresponds to one unique orbit $\mathcal{O}(i,j,k,l)$. After Algorithm 1 terminates, $\mathscr{S}_{:,:,:,:,m}$ (letting the first four indices of $\mathscr{S}$ be free but fixing the last index at $m$) will be the $m$-th element of the basis, an $N\times N\times N\times N$ tensor that satisfies the symmetry (17). In the algorithm, the purpose of the binary tensor $\mathscr{B}$ is to signal whether we have already captured the orbit $\mathcal{O}(i,j,k,l)$. Upon the algorithm’s termination, $\mathscr{B}$ will be identically equal to $1$ and $m$ will equal $n_{T}$. With this framework, we build the model
$\tau_{ijkl}$ $=\sum_{m=1}^{n_{T}}\mathscr{S}_{ijklm}\beta_{m}$ (19a)
$\widetilde{H}(P)_{ij}$ $=H^{\text{core}}_{ij}+\sum_{kl}\left[\tau_{ijlk}-\frac{1}{2}\tau_{iklj}\right]P_{kl}.$ (19b)
As $\tau$ is a linear combination of basis elements from $\mathscr{S}$, it is clear that $\tau$ will satisfy the $8$-fold symmetry (17). In this model, as long as $P$ is Hermitian, then $H$ will be Hermitian too:
$\overline{\widetilde{H}(P)_{ji}}=\overline{H^{\text{core}}_{ji}}+\sum_{kl}\left[\tau_{jilk}-\frac{1}{2}\tau_{jkli}\right]\overline{P_{kl}}\\ =H^{\text{core}}_{ij}+\sum_{lk}\left[\tau_{ijkl}-\frac{1}{2}\tau_{ilkj}\right]P_{lk}=\widetilde{H}(P)_{ij}.$
For the model (19), the parameters consist of $\bm{\beta}\in\mathbb{R}^{n_{T}}$. This is roughly $1/8$-th as many parameters as needed in the models (11) and (16).
In SI Section V.2.3, we explain how to compute $\bm{\beta}$ such that the model (19) matches the ground truth Hamiltonian.
II.3 Generation of data sets for training and testing
To train and test our models, we follow procedures for data generation, taking care to keep these procedures uniform across all molecular systems that we study. In what follows, $H(P)$ can either be the ground truth Hamiltonian, or it can be one of the $\widetilde{H}(P)$ models described above.
Single field-free trajectory: As described in the first three paragraphs of Section II.1 above, we take as our initial condition an electron density matrix $P(0)$ that has been advanced slightly past the time of an initial delta-kick perturbation. We choose the delta-kick perturbation because it is maximally delocalized in Fourier space (equivalently, maximally localized in time), leading to trajectories that explore more modes of oscillation, which in turn helps our learning procedures to reconstruct the global inter-electronic potential. Using this $P(0)$ and a field-free Hamiltonian $H(P)$, we apply the CI4 method to numerically solve (2) and compute a single field-free trajectory $\{P(t_{j})\}_{j=0}^{J}$. Here and in what follows, we use the temporal grid $t_{j}=j\Delta t$ and $J=200000$ integration steps. Unless specified otherwise, $\Delta t=8.268\times 10^{-4}$ a.u.
When we use this trajectory for training, we also compute time-derivatives $\{\dot{P}(t_{j})\}_{j=2}^{J-2}$ using the fourth-order centered-difference formula
$\dot{P}(t_{j})=\frac{-P(t_{j+2})+8P(t_{j+1})-8P(t_{j-1})+P(t_{j-2})}{12\Delta t}+O(\Delta t^{4}).$ (20)
We then restrict the trajectory to match the points in time at which we have time-derivatives, i.e., we pair $\{\dot{P}(t_{j})\}_{j=2}^{J-2}$ with $\{P(t_{j})\}_{j=2}^{J-2}$ and use the pair for training.
Ensemble of field-free trajectories: We take the $P(0)$ described in the previous paragraph and perturb it using a random Hermitian matrix $\mathscr{H}$, creating a new initial condition $\mathscr{P}(0)=P(0)+\varepsilon\mathscr{H}$. To generate $\mathscr{H}$, we start with $2N^{2}$ independent samples from the standard normal distribution, from which we form two real matrices $\mathscr{D}_{R}$ and $\mathscr{D}_{I}$ both of size $N\times N$. Let denote conjugate transpose. We take $\mathscr{D}=\mathscr{D}_{R}+\imath\mathscr{D}_{I}$ and then set $\mathscr{H}=(\mathscr{D}+\mathscr{D}^{\dagger})/2$. In this paper, we set $\varepsilon=(10/N^{2})\sum_{i,j}|P_{ij}(0)|$, i.e., $10$ times the mean absolute value of $P(0)$. Though $\mathscr{P}(0)$ is Hermitian, we must modify it slightly to be idempotent, a requirement for Hartree-Fock electron density matrices. We compute the eigendecomposition $\mathscr{P}(0)=VDV^{\dagger}$. Any entries of $D$ greater than $1/2$ are set to $1$, while entries at most equal to $1/2$ are set to 00. With this modified $D$, the resulting $\mathscr{P}(0)=VDV^{\dagger}$ is guaranteed to be idempotent.
Using the above procedure, we generate a cloud of Hermitian, idempotent initial conditions $\{\mathscr{P}_{k}(0)\}_{k=1}^{100}$. Given a field-free Hamiltonian $H(P)$ and this cloud of initial conditions, we use CI4 (with $\Delta t=8.268\times 10^{-4}$ a.u.) to numerically integrate forward in time for $J=20000$ steps. With $t_{j}=j\Delta t$, this yields a cloud of trajectories $\{\mathscr{P}_{k}(t_{j})\}_{k=1}^{100}$ for $j=0,\ldots,J$. We apply (20) to compute time-derivatives $\{\dot{\mathscr{P}}_{k}(t_{j})\}_{k=1}^{100}$ on the equispaced grid $t_{j}=j\Delta t$ for $j=2,\ldots,J-2$. We restrict $\{\mathscr{P}_{k}(t_{j})\}_{k=1}^{100}$ to the same temporal grid, so that trajectories and their time-derivatives are matched.
[…]
When we numerically solve (2) in any setting, we monitor $P(t)$ at every time step to ensure that it remains Hermitian and idempotent.
II.4 Training
For each $\widetilde{H}$ model we have proposed, let $\bm{\theta}$ denote the vector of model parameters. For the tied regression model (11) and the $8$-fold symmetry-preserving model (19), $\bm{\theta}=\bm{\beta}$. For the Hermitian representation model (16), $\bm{\theta}=(\mathbf{v},\mathbf{w})$. In this section, we will emphasize the parametric dependence of each $\widetilde{H}$ model on its parameters $\bm{\theta}$ by writing $\widetilde{H}(P;\bm{\theta})$.
To train a given $\widetilde{H}(P;\bm{\theta})$ model, we need training data. For a given molecular system (choice of molecule plus basis set), we use the true Hamiltonian (8) in the field-free procedures described in Section II.3. This yields two different training sets (single field-free trajectory and ensemble of field-free trajectories). Both training sets are of the form $\{(P^{(j)},\dot{P}^{(j)})\}_{j=1}^{n_{\text{train}}}$; each pair in this set consists of an electron density matrix and its time-derivative, taken from the same trajectory and evaluated at the same time.
Given one of the three models (11), (16), or (19), and either training set, the residual associated with the $j$-th data point is the difference between the left- and right-hand sides of the TDHF equation (2) with the time-derivative $dP/dt$ approximated via $\dot{P}$ defined in (20):
$S^{(j)}(\bm{\theta})=\imath\dot{P}^{(j)}-\big{[}\widetilde{H}(P^{(j)};\bm{\theta}),P^{(j)}\big{]}$ (21)
We then take as our loss function the sum of squared magnitudes of all residuals:
$\mathcal{L}(\bm{\theta})=\sum_{j=1}^{n_{\text{train}}}\sum_{a,b}|S^{(j)}_{ab}(\bm{\theta})|^{2}.$ (22)
If we were to use the ground truth Hamiltonian $H(P)$ instead of a model $\widetilde{H}(P)$, the loss would be practically zero. Therefore, to train the model, we seek $\bm{\theta}^{\star}=\min_{\bm{\theta}}\mathcal{L}(\bm{\theta})$. As each of our models $\widetilde{H}(P)$ is linear in the parameters $\bm{\theta}$, so is the residual (21). Hence the loss (22) is quadratic in $\bm{\theta}$; quadratic minimization problems have been well-studied. To leverage known techniques, we put our problem into standard form. As the residual is linear in the parameters $\bm{\theta}$, we can write it as
$S^{(j)}(\bm{\theta})=\underbrace{\begin{bmatrix}\nabla_{\bm{\theta}}S^{(j)}\end{bmatrix}}_{A^{j}}\bm{\theta}+\underbrace{S^{(j)}(\mathbf{0})}_{-\mathbf{c}^{j}}$ (23)
Let $|\bm{\theta}|$ denote the length of the vector $\bm{\theta}$. In (23), we interpret $S^{(j)}$ as a flattened vector in $\mathbb{C}^{N^{2}}$, so that the Jacobian $A^{j}=\nabla_{\bm{\theta}}S^{(j)}$ is a matrix of shape $N^{2}\times|\bm{\theta}|$. We treat $\mathbf{c}^{j}$ as a column vector of shape $N^{2}\times 1$. Stacking objects vertically, we can rewrite the loss (22) as
$\mathcal{L}(\bm{\theta})=\left\|\begin{bmatrix}A^{1}\\ \vdots\\ A^{n_{\text{train}}}\end{bmatrix}\bm{\theta}-\begin{bmatrix}\mathbf{c}^{1}\\ \vdots\\ \mathbf{c}^{n_{\text{train}}}\end{bmatrix}\right\|^{2}=\|\mathbf{A}\bm{\theta}-\mathbf{C}\|^{2}.$ (24)
[…]
II.4.1 Gradient- and Hessian-based minimization of the loss
If we take the gradient of the loss (24) with respect to $\bm{\theta}$ and set it equal to zero, the resulting solution can be expressed in terms of (i) the Hessian of the loss and (ii) the gradient of the loss at $\bm{\theta}=\mathbf{0}$:
$\bm{\theta}^{\star}$ $=-(\mathbf{Hess}\ \mathcal{L})^{-1}\left[\nabla_{\bm{\theta}}\mathcal{L}(\mathbf{0})\right]$ (25a)
[…]
As $\mathbf{Hess}$ is of size $|\bm{\theta}|\times|\bm{\theta}|$, we can only use (25a) to train our models when the number of parameters $|\bm{\theta}|$ is sufficiently small. When $\mathbf{Hess}$ has eigenvalues near zero, we replace the matrix inverse with the Moore-Penrose pseudoinverse (Hansen, 1998) with threshold $10^{-12}$. Here we treat problems for which $N$ can be as large as $60$; in this case, for the models (11) and (16), $|\bm{\theta}|=N^{4}=12.96\times 10^{6}$. Assuming $8$-byte double precision floats, a Hessian of size $N^{4}\times N^{4}$ would consume over $1$ petabyte, motivating the use of matrix-free methods.
II.4.2 Matrix-free minimization of the loss
Returning to the loss (24), we see that the critical objects are the Jacobians $A^{j}=\nabla_{\bm{\theta}}S^{(j)}$, which can be written as $A^{j}_{k,l,m}={\partial S^{(j)}_{kl}}/{\partial\theta_{m}}$. Thus $\mathbf{A}$ is a $4$-index tensor of dimension $n_{\text{train}}\times N\times N\times|\bm{\theta}|$. Since we will not (in general) be able to store $\mathbf{A}$ in memory, we will present efficient numerical algorithms to compute the left and right products of $\mathbf{A}$ against tensors of appropriate sizes. We will use these algorithms as inputs to the matrix-free method LSMR (Fong and Saunders, 2011), and thereby compute minimizers of the loss (24).
Note that in situations where $\mathbf{A}$ does not have full column rank, LSMR returns the solution $\bm{\theta}^{\star}$ with minimal $2$-norm; this is particularly useful for problems such as ours, where $\mathbf{Hess}$ may have nearly zero eigenvalues and high condition number. Our prior work combatted these problems by combining Hessian-based training with a Tikhonov or ridge regularization approach (Gupta et al., 2022). For the present work, when we tried adding such regularization terms to the loss, we found that results did not improve, likely due to the implicit regularization already present in LSMR.
Please see SI Section V.3 for detailed expressions of Jacobian-vector products (both left and right) for all three models (11), (16), and (19). In SI, we also explain how we implement these products efficiently, so that we can train our models even when they contain tens of millions of parameters. For model (19), we also provide (in SI Section V.3.3) expressions for the gradient and Hessian of the loss $\mathcal{L}$, which are used to train model (19) for small molecular systems.
III Results
We test our models on seven molecular systems, each consisting of a molecule together with a particular atom-centered Gaussian basis set: (i) $\text{HeH}^{+}$ with 6-31G ($N=4$), (ii) $\text{HeH}^{+}$ with 6-311G ($N=14$), (iii) LiH with 6-31G ($N=11$), (iv) $\text{C}_{2}\text{H}_{4}$ with STO-3G ($N=14$), (v) LiH with 6-311++$\text{G}^{**}$ ($N=29$), (vi) $\text{C}_{2}\text{H}_{4}$ with 6-31+$\text{G}^{*}$ ($N=46$) and (vii) $\text{C}_{6}\text{H}_{6}\text{N}_{2}\text{O}_{2}$ with STO-3G ($N=60$). In SI Section V.5, we provide the Cartesian coordinates for each molecule.
III.1 Model training
For models (11) and (16), the $N^{4}$ scaling of the number of parameters leads us to use the iterative, matrix-free LSMR method to train. For the $8$-fold symmetry-preserving model (19), we train using LSMR for all molecular systems—though this is only necessary when $N\geq 29$, i.e., for molecular systems (v), (vi), and (vii). Molecular systems (i), (ii), (iii), and (iv) are sufficiently small such that we can train model (19) using the exact Hessian-based formula (25a). We carry this out for the purposes of comparison.
In turn, we train all models on the single field-free trajectory and on the ensemble of field-free trajectories described in Section II.3. There are small differences in how we treat the smaller molecular systems (i-iv) versus the larger systems (v-vii). Let us first describe these differences for training on single trajectory data. For the smaller systems, when we use LSMR, we use the entire matched data set $\{P(t_{j}),\dot{P}(t_{j})\}_{j=2}^{j=J-2}$ with $J=200000$. For the larger systems, we take every $10$th snapshot in time, so that our training set is of size $20000$.
The ensemble training set initially consists of a cloud of $100$ matched trajectories $\{\mathscr{P}_{k}(t_{j}),\dot{\mathscr{P}}_{k}(t_{j})\}_{k=1}^{100}$ for $j=2,\ldots,J-2$ with $J=20000$. For the smaller molecular systems, we take every $50$th snapshot from this cloud, for a total of $40000$ temporal snapshots ($400$ snapshots from each of the $100$ trajectories). We augment this using $40000$ snapshots from the single field-free trajectory and its time-derivative, i.e., we augment using every $5$th snapshot from $\{P(t_{j}),\dot{P}(t_{j})\}_{j=2}^{j=199998}$. In total, for smaller molecular systems, our ensemble training set consists of $80000$ snapshots. For the larger molecular systems, we take every $100$ snapshot from the cloud and augment with every $10$th snapshot from the single field-free trajectory, leading to a total ensemble training set size of $40000$ snapshots.
We use Tied, Herm, and Symm to denote the respective $\widetilde{H}(P)$ models (11), (16), and (19), each trained on the single field-free trajectory using the matrix-free LSMR algorithm. We use TiedE, HermE, and SymmE to denote the respective $\widetilde{H}(P)$ models (11), (16), and (19), each trained on the ensemble of field-free trajectories using the matrix-free LSMR algorithm. Finally, we use SymmH and SymmHE to denote the $8$-fold symmetry-preserving $\widetilde{H}(P)$ model (19), trained via the exact Hessian-based formula (25a), on the single trajectory or ensemble training sets, respectively.
In summary, we obtain $8$ trained models for each of the $4$ smaller molecular systems (i), (ii), (iii), and (v). We obtain $6$ trained models for each of the $3$ larger molecular systems (iv), (vi), and (vii).
(a) $\text{HeH}^{+}$ in 6-31G
(b) $\text{HeH}^{+}$ in 6-311G
(c) LiH in 6-31G
(d) $\text{C}_{2}\text{H}_{4}$ in 6-31G
(e) LiH in 6-311++$\text{G}^{**}$
(f) $\text{C}_{2}\text{H}_{4}$ with 6-31+$\text{G}^{*}$
(g) $\text{C}_{6}\text{H}_{6}\text{N}_{2}\text{O}_{2}$ in STO-3G
Figure 2: We plot the training loss versus iteration number for six models and seven molecular systems. Here Tied, Herm, and Symm refer to the models (11), (16) and (19), respectively. The suffix “E” indicates that the model has been trained on an ensemble of field-free trajectories; otherwise, the model has been trained on a single field-free trajectory. Overall, we see that the $8$-fold symmetry-preserving model with the ensemble training set (SymmE) achieves the best results, and is significantly better for the larger molecular systems (e-g). For the four smaller molecular systems (a-d), ensemble training reduces the number of iterations by orders of magnitude for all models.
III.2 Metrics
In Figure 2, for the models trained with the iterative LSMR method, we plot training loss as a function of the number of iterations. Across all molecular systems, the SymmE model achieves the best results. For larger molecular systems, the SymmE model requires orders of magnitude fewer number of iterations to achieve small values of the training loss. For smaller molecular systems, any of the models trained on the ensemble of field-free trajectories (TiedE, HermE, and SymmE) yield the best performance in terms of rapid minimization of training loss, i.e., within $10^{3}$ iterations. For a detailed table and plot of final values of the training loss for all models and all molecular systems, see SI Section V.4.1.
After training, we assess model errors using three metrics: propagation error, Hamiltonian error, and commutator error.
(a) Field-free propagation errors
(b) Field-on propagation errors
Figure 3: For the four smaller molecular systems, we plot $\infty$-norm propagation errors (27) for the $8$-fold symmetry-preserving model (19) trained using LSMR (Symm, SymmE) and also using the exact Hessian-based approach (SymmH, SymmHE) in (25a). The suffix “E” indicates that the model has been trained on an ensemble of field-free trajectories; otherwise, the model has been trained on a single field-free trajectory. The results show that the iterative LSMR method leads to models that compare very favorably against those trained via the exact Hessian-based approach. This is true for both field-free (training, left) and field-on (test, right) propagation errors.
III.2.1 Propagation Error
Propagation error measures the ability of the learned Hamiltonian $\widetilde{H}(P)$ to generate predicted trajectories $\widetilde{P}(t)$ that match true trajectories $P(t)$. Let $P(t)$ denote the true trajectory obtained by solving (2) using the true Hamiltonian $H(P)$, and let $\widetilde{P}(t)$ denote the solution of the TDHF equation
$\imath\frac{d\widetilde{P}(t)}{dt}=\big{[}\widetilde{H}(t,\widetilde{P}(t)),\widetilde{P}(t)\big{]},$ (26)
with a Hamiltonian $\widetilde{H}(t,P)$ that equals the learned field-free Hamiltonian $\widetilde{H}(P)$ plus an optional time-dependent term. As explained in Section II, we use the CI4 scheme to solve both (2) and (26) for $J=20000$ time steps at $\Delta t=8.268\times 10^{-4}$, starting both trajectories at the same initial condition $\widetilde{P}(0)=P(0)$. We then compute the $\infty$-norm error
$\|P-\widetilde{P}\|_{\infty}=\max_{1\leq j\leq J}\max_{1\leq a,b\leq N}|P(t_{j})_{ab}-\widetilde{P}(t_{j})_{ab}|,$ (27)
the worst-case error across all matrix entries and all time.
We compute propagation error in both field-free and field-on settings. In the field-free setting, we solve (2) and (26) using the true and learned field-free Hamiltonians, respectively. We view the resulting field-free propagation errors as training set errors. In the field-on setting, we add a time-dependent electric field term to both the true and learned field-free Hamiltonians—see Section II.3. As the true field-on trajectory is not used for training, we view the resulting field-on propagation errors as test set errors.
As mentioned in Section II.2.3, the reduced number of parameters in the $8$-fold symmetry-preserving method enables use of the exact Hessian-based formula (25a) to train models for the four smaller molecular systems. Thus we can use these four smaller systems to study how much error (if any) we introduce by training the $8$-fold model via the iterative LSMR method versus the exact Hessian-based approach.
In Figure 3, we show that in terms of propagation error (27) on both the field-free/training set (left) and the field-on/test set (right), the LSMR-trained models do no worse—and in some cases outperform—the Hessian-trained models. These results justify our use of the LSMR method to train the larger molecular systems, for which a comparison against the Hessian-based approach is impossible. For the remainder of this paper, we focus our attention on LSMR-trained models. Further results for Hessian-trained models can be found in SI Section V.4.2.
(a) Field-free propagation errors
(b) Field-on propagation errors
Figure 4: For all seven molecular systems, we plot $\infty$-norm propagation errors (27) for all six LSMR-trained models in both field-free/training (left) and field-on/test (right) settings. Overall, the $8$-fold symmetry-preserving model SymmE, trained on an ensemble of trajectories, is either the best (e.g., for $\text{C}_{6}\text{H}_{6}\text{N}_{2}\text{O}_{2},$ the largest system) or incurs error that only slightly exceeds that of the best model.
In Figure 4, we plot the propagation error (27) for all LSMR-trained models and all seven molecular systems, on both the field-free/training set (left) and the field-on/test set (right). All reported errors are orders of magnitude lower than what we achieved in prior results (Bhat, Ranka, and Isborn, 2020; Gupta et al., 2022). In nearly all cases, and especially for the three largest systems, ensemble-trained models (TiedE, HermE, and SymmE) outperform single trajectory-trained models (Tied, Herm, and Symm), often by several orders of magnitude. For the two small systems $\text{HeH}^{+}$ and LiH in 6-31G, the only model that outperforms ensemble-trained models is Symm, the $8$-fold symmetry-preserving model trained on a single field-free trajectory.
Among all models, the $8$-fold symmetry-preserving model SymmE, trained on an ensemble of trajectories, is nearly the best (if not the best) for all molecular systems. For $\text{C}_{6}\text{H}_{6}\text{N}_{2}\text{O}_{2},$ the largest system we considered, the difference between SymmE and other models is especially striking.
As long as $\Delta t>0$, the training loss (22) measures something different than propagation error on the training set. When we compute the loss, we are measuring our ability to predict a discrete-time version of the derivative $\dot{P}(t_{j})$ using the true value of $P(t_{j})$ at each time. This is equivalent to predicting a small number of time steps (at most, the width of the finite-difference stencil used to compute $\dot{P}$) ahead in time. Thus the loss measures aggregate short-term propagation error. When we compute the propagation error metrics (27) and (28), we use the predicted trajectory $\widetilde{P}(t_{j})$, which can potentially accumulate errors at each time step. There is no $t_{j}>0$ at which we reset $\widetilde{P}(t_{j})$ to equal $P(t_{j})$. Thus the metrics (27) and (28) measure aggregate long-term propagation error.
In SI Section V.4.2, we plot—for each model and each molecular system—the time-dependent mean absolute error (MAE)
$\text{MAE}(t_{j})=\frac{1}{N^{2}}\sum_{a,b}|P(t_{j})_{ab}-\widetilde{P}(t_{j})_{ab}|,$ (28)
a scalar time series showing how the error evolves in time. We use MAE as it conveys the average number of digits of accuracy in each entry of a predicted electron density matrix. The MAE results agree with the $\infty$-norm results in Figure 4.
Figure 5: We plot Hamiltonian and commutator errors for all molecular systems and all LSMR-trained models. The left plot shows that SymmE, the $8$-fold symmetry-preserving model (19) trained on the ensemble of field-free trajectories, has the lowest Hamiltonian error (29) for each molecular system. The middle and right plots show $\infty$-norm commutator error (30) in the field-free/training and field-on/test cases, respectively. The commutator errors reveal that all ensemble-trained models perform similarly well, except for $\text{C}_{6}\text{H}_{6}\text{N}_{2}\text{O}_{2},$ for which SymmE shows a large advantage over all other models.
III.2.2 Hamiltonian and Commutator Errors
For each of the three models we have proposed, there exists a set of parameters $\bm{\theta}^{\circ}$ such that the model perfectly matches the ground-truth Hamiltonian (8)—for detailed derivations of this, see SI Section V.2. Thus it makes sense to quantify, for each model, how far the trained parameter values $\bm{\theta}^{\star}$ are from the ground truth parameters $\bm{\theta}^{\circ}$. For the three respective models (11), (16), and (19), we compute the $\infty$-norm error
$\|\bm{\theta}^{\star}-\bm{\theta}^{\circ}\|_{\infty}=\max_{1\leq j\leq|\bm{\theta}|}|\theta^{\star}_{j}-\theta^{\circ}_{j}|.$ (29)
For each model, we view the parameters as a flattened vector $\bm{\theta}$ of length $|\bm{\theta}|$. We use $\bm{\theta}^{\ast}$ and $\bm{\theta}^{\circ}$ to denote, respectively, the trained and true values of the parameters.
In the left-most plot in Figure 5, we find that SymmE, the $8$-fold symmetry-preserving model (19) trained on the ensemble of field-free trajectories, has the smallest Hamiltonian error (29) across all molecular systems and all LSMR-trained models.
Examining the equations of motion (2), we see the possibility that some part of the Hamiltonian may not affect the dynamics at all. Specifically, if $\widetilde{H}=\widetilde{H}_{\text{comm}}+\widetilde{H}_{\text{rem}}$ where $\widetilde{H}_{\text{comm}}$ commutes with $P$, then $[\widetilde{H},P]=[\widetilde{H}_{\text{rem}},P]$ where $\widetilde{H}_{\text{rem}}$ stands for the remaining part of $\widetilde{H}$. If the errors in the learned Hamiltonian $\widetilde{H}$ concentrate in $\widetilde{H}_{\text{comm}}$, then those errors will cancel out of the commutator. To quantify the extent to which this happens, we compute commutator errors in the $\infty$-norm:
$\|[H(P(t),t),P(t)]-[\widetilde{H}(P(t),t),P(t)]\|_{\infty}\\ =\|[H(P(t),t)-\widetilde{H}(P(t),t),P(t)]\|_{\infty}$ (30)
The Hamiltonian only appears in the equation of motion (2) through the commutator. If the commutator error is small, the learned Hamiltonian $\widetilde{H}$ is accurate for practical purposes.
In the middle (field-free/training) and right-most (field-on/test) plots of Figure 5, we plot the commutator error for all molecular systems and all LSMR-trained models. We see that the TiedE and HermE models have nearly the same commutator error as the SymmE model, for all molecular systems except for $\text{C}_{6}\text{H}_{6}\text{N}_{2}\text{O}_{2}$ —here SymmE maintains its superior performance. These results explain why the TiedE, HermE, and SymmE models all have low and nearly equal propagation errors (see Figure 4), even though the raw Hamiltonian error of the TiedE and HermE models can be large (left-most plot of Figure 5).
Finally, note that the field-free and field-on commutator errors show extremely similar trends. This is because, in the field-on case, the electric field terms in $H$ and $\widetilde{H}$ cancel perfectly, so that the commutator error reduces to $\|[H(P(t))-\widetilde{H}(P(t)),P(t)]\|$ where $H$ and $\widetilde{H}$ are the field-free Hamiltonians and $P(t)$ is the field-on trajectory.
Further tables of results can be found in SI Section V.4.3.
IV Discussion and Conclusion
As emphasized in Section I, we undertook this study to develop strategies to learn quantum potentials from time-dependent data. We focused on time-dependent Hartree-Fock (TDHF) systems because we know the full form of the field-free Hamiltonian, we have no problem using TDHF to generate data for large molecular systems, and because the lessons we learn here can be cross-applied to more complex approaches such as time-dependent density functional theory (TDDFT), time-dependent reduced density matrix functional theory (TDRDMFT), or Kadanoff-Baym theory.
One clear conclusion is the importance of both the quality and quantity of the training data. For the largest molecular systems under study, we were able to achieve acceptably low values of field-on (test set) propagation error only by training on data generated via the fourth-order CI4 scheme with a sufficiently small time step ($\Delta t$). With the coarser second-order MMUT scheme and/or a larger time step, we would not have been able to achieve these results.
In our prior work, we always trained our models on single field-free trajectories (Bhat, Ranka, and Isborn, 2020; Gupta et al., 2022). The results presented in Section III show the importance of training with an ensemble of field-free trajectories. In Section II.3, we described the construction of this ensemble in detail. In particular, when we constructed random initial conditions for the cloud of trajectories in the ensemble, we ensured that these initial conditions were Hermitian and idempotent. Indeed, if we train on ensembles generated from trajectories that ignore these constraints, the resulting models perform poorly. Especially for larger molecular systems, training on physically meaningful ensembles constrains our models to stay closer to the true Hamiltonian than is possible with a single trajectory training set.
We explored three models in this work, two of which preserve the Hermitian symmetry of the Hamiltonian, and the third of which preserves a deeper $8$-fold symmetry. Taking a holistic view of the training loss versus number of iterations, the propagation error, the Hamiltonian error, and the commutator error, we can build a strong case that the best model is the $8$-fold symmetry-preserving, ensemble-trained model (SymmE). In short, building more symmetries (i.e., more physics) into the machine learning model yields better results, especially as molecular system size increases.
We have focused on a method (TDHF) in which the underlying symmetries are known exactly. In cases where we seek to learn a term (e.g., an exchange-correlation potential or self-energy) whose full symmetries are unknown, the present paper offers two approaches. First, if there is a symmetry that one suspects should be present, one can easily generalize our methods to create an appropriately symmetry-constrained machine learning model. For a linear model, the only substantive changes to Algorithm 1 will concern the dimensions of the tensors and the elements of the symmetry group $\{\sigma^{m}\}$. For nonlinear neural network models, one can use or adapt methods explored previously (Anderson, Hy, and Kondor, 2019; Pan and Kondor, 2022). If the resulting model yields results that are significantly better than non-symmetrized models, then a reasonable working hypothesis is that the unknown term truly possesses the enforced symmetry.
Second, our results show that preserving only the Hermitian symmetry leads to predictive accuracy in terms of long-term propagation error: the TiedE and HermE models perform well in propagation error and commutator error tests. Though they cannot compare with SymmE in terms of ease of training (training loss versus iteration number) or $\infty$-norm Hamiltonian error, the TiedE and HermE models were far easier to implement (in terms of code development time) than the SymmE model. Generalizing to more complex situations such as TDDFT or TDRDMFT, we hypothesize that a machine learning model that preserves only an outer layer of symmetries may be sufficient for a first round of results. In subsequent rounds, these models should be refined by incorporating more symmetries, especially if the goal is to learn the true Hamiltonian. This will ensure that the model can extrapolate beyond the training data, and also increase the chances that one can distill new physical insights from, e.g., a machine-learned closure model for theories such as TDDFT and TDRDMFT.
Acknowledgements.
This research was sponsored by the Office of Naval Research and was accomplished under Grant Number W911NF-23-1-0153. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. This research was also supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences under Award Number DE-SC0020203. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231 NERSC award ASCR-ERCAP0028496. We acknowledge computational time on the Pinnacles cluster at UC Merced (supported by NSF OAC-2019144). We also acknowledge use of resources available through the National Research Platform (NRP) at the University of California, San Diego. NRP has been developed, and is supported in part, by funding from the National Science Foundation, from awards 1730158, 1540112, 1541349, 1826967, 2112167, 2100237, and 2120019, as well as additional funding from community partners.
Data Availability Statement
The data that support the findings of this study are openly available in the GitHub repository learningTDHF at
https://github.com/hbhat4000/learningTDHF/.